OpenStudy (chillout):

A RC circuit containing a battery, a capacitor and a resistor: Given that R*dQ/dt + Q/C = V Solve the differential equation for Q(t). A little help here?

OpenStudy (anonymous):

\[R * dQ/dt +Q/ C = V \] You may or may not recognize this as a first order, linear differential equation, http://www.sosmath.com/diffeq/first/lineareq/lineareq.html

OpenStudy (chillout):

Gonna try to solve it, one moment.

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

Yeah One Moment Up d00d, Your time is Up ! :P :D

OpenStudy (chillout):

I've just started studying ODEs, don't be hard on me :(

OpenStudy (anonymous):

Hey I was pulling your leg d00d, try it. Soon you become expert at it! Good luck!

OpenStudy (chillout):

All right, here's what I did. I separated the variables and got\[\frac{ R*dQ }{ dt } = V - \frac{ Q }{ C } \leftarrow \rightarrow \frac{ dQ }{ V - C } = \frac{ dt }{ R }\] Integrating both sides I got\[\ln |V-\frac{ Q }{ C }| = \frac{ t }{ -c*R } + c\] I guess it's pretty easy noe :P

OpenStudy (anonymous):

How did u get that?

OpenStudy (chillout):

Just Ignore that "-c"

OpenStudy (anonymous):

\[ \frac{ R*dQ }{ dt } +\frac{ Q }{C }=V\] \[ \frac{ R*dQ }{ dt }= V-\frac{ Q }{C }\] \[ \frac{ R*dQ }{ dt }= \frac{ VC-Q }{C }\] \[ \frac{ dQ }{ VC-Q }= \frac{ dt }{CR }\]

OpenStudy (chillout):

Oh! Noticed what went wrong... damn, pretty dumbish error lol

OpenStudy (anonymous):

Hmm! it happens sometimes, never mind it!

OpenStudy (chillout):

Thanks for the help! I've reached the answer.

OpenStudy (anonymous):

If you are really interested in learning ODE's I suggest this video http://www.youtube.com/watch?v=CogfMjKUGc0 [ the video also contains some programming in between, never mind it , just concentrate on theory he is saying]

OpenStudy (chillout):

That will help, thanks.

OpenStudy (anonymous):

My pleasure But be patient while hearing that lecture, coz that lecture might be too long for you!